OpenStudy (anonymous):

Solve for C -14=-5+3c

OpenStudy (anonymous):

-5+3=-2 14/(divided by)-2=-7 c=-7

OpenStudy (anonymous):

c=3 trust me

OpenStudy (anonymous):

C=7* not -7 sorry

OpenStudy (anonymous):

yes i am right

OpenStudy (anonymous):

Tony, you're wrong -5+3=-2 -2x3= -6, not -14